# Co-Evolving Agents: Learning from Failures as Hard Negative

### v1

@@ Abstract
The rapid progress of large foundation models has accelerated the development of task-specialized agents across diverse domains. However, the effectiveness of agents remains tightly coupled with the quality of training data, while curating task-specific datasets remains costly and often infeasible in real-world scenarios. Recent work has explored self-improving agents that autonomously generate, refine, and re-train on their own trajectories. A prominent line of approaches further leverages preference optimization by pairing predicted trajectories with scarce ground-truth trajectories, enabling agents to learn directly from their own failures. While these methods outperform supervised fine-tuning, their heavy reliance on predicted trajectories under limited ground-truth supervision leaves them prone to overfitting. To address this, we propose a co-evolving agents framework in which a target agent improves jointly with an auxiliary failure agent. The failure agent learns through preference optimization over failure trajectories from both the target and itself, thereby generating hard negatives that are close to success yet remain failures. Incorporating these informative hard negatives into the target agent’s optimization sharpens decision boundaries and enhances generalization. Our comprehensive analysis and experiments across benchmark datasets show that our method not only show improved performance but also highlights that failures, instead of being used as-is, can be systematically transformed into structured and valuable learning signals in self-improving agents.
@@ 1 Introduction
The rapid progress of large foundation models (gpt5; qwen3; llama4; claude; gemini) has facilitated the rise of task-specialized agents across diverse domains, ranging from open-domain dialogue to scientific reasoning tasks (learnbyinteract; agenttuning; agentrefine; robocat). These agents inherit the broad generalization capacity of pretrained models, allowing effective adaptation to new tasks with relatively limited supervision. This promise has motivated growing interest in developing methods that adapt foundation models into reliable and effective domain-specialized agents. Recent advances in multi-agent systems and preference optimization further highlight the potential of combining broad pretraining with specialized adaptation. In particular, the ability to automatically curate training signals from agent interactions opens opportunities for scaling beyond static, human-labeled corpora. At the same time, the increasing deployment of agents in dynamic, real-world settings emphasizes the importance of approaches that can continuously refine behavior without costly retraining. Such developments make it timely to revisit how failures, long viewed as undesirable artifacts, can instead be leveraged as constructive learning signals.
@@
Nevertheless, the effectiveness of such agents remains constrained by the quality of task-specific training data (star; clues). High-quality datasets are essential for reliable reasoning and decision making, providing the signals required for adaptation to specialized domains. Yet, constructing such datasets is expensive and labor-intensive, often requiring domain expertise and extensive annotation. In many real-world scenarios, constructing large curated datasets is infeasible. In addition to the prohibitive cost of collecting interactions at scale, the need to repeatedly curate data to keep pace with non-stationary environments makes this approach impractical. This bottleneck has motivated growing interest in methods that enable agents to improve autonomously without relying on continuous manual curation (agentR; dynasaur; golden). Such methods aim to bridge the gap between the flexibility of foundation models and the rigorous demands of domain-specific tasks, while keeping human intervention to a minimum. A central challenge is therefore to design mechanisms that transform abundant but noisy interaction data into structured supervision that drives reliable improvement.
@@
Self-improving agents (learnbyinteract; agenttuning; agentrefine) have emerged as a promising paradigm to reduce reliance on costly human annotation. Maintaining agents at state-of-the-art performance would require continuous human annotation, which is prohibitively costly and infeasible at scale. Instead, self-improving agents automate parts of the data construction process by synthesizing expert-like trajectories from external resources such as documentation or databases, and by repurposing predicted failures as preference data for training. However, in challenging downstream tasks where pretrained LLMs perform poorly, generating high-quality trajectories themselves remains a major bottleneck, making fully autonomous self-improvement difficult in practice.
@@
In this context, Exploration-Based Trajectory Optimization (ETO) (eto) constructs preference datasets by pairing failure trajectories predicted by the agent with ground-truth ones using a given reward model, enabling preference optimization (dpo). Despite the promise of autonomous improvement, this approach remains limited, as it relies on the model’s own failure trajectories. These failures are substantially less informative than human-curated negatives, especially in tasks where pretrained LLMs lack prior knowledge. As a result, the dispreferred trajectories offer only weak contrast, making DPO focus on simply increasing the likelihood of expert trajectories rather than shaping a fine-grained decision boundary. This weak supervision often yields limited corrective signals, causing the model to gravitate toward the small expert set and ultimately overfit, rather than acquiring a robust understanding of the underlying task landscape.
@@
To overcome these limitations, we introduce a co-evolving agents framework in which a target agent improves jointly with an auxiliary failure agent. The failure agent specializes in preference optimization over failure trajectories from both the target and itself, enabling it to learn a fine-grained landscape of failures rather than merely preferring expert trajectories. Crucially, this design enables the agent to autonomously generate informative hard negatives (hardnegative; dpo; simclr) that are high-reward failures close to success, without any external supervision. Incorporating these hard negatives into the target agent’s preference optimization provides stronger and more diverse contrastive signals, sharpening decision boundaries and yielding more generalizable performance.
We validate our framework through comprehensive analysis and experiments across diverse domains, including the online shopping environment WebShop (webshop), the science reasoning environment ScienceWorld (scienceworld), and the interactive SQL environment InterCodeSQL (intercode). Our analysis on these benchmarks verifies that the failure agent does not simply imitate expert trajectories but continues to generate high-reward failures that serve as informative hard negatives.
Experiments further demonstrate substantial improvements over competitive baselines, achieving large margins of gain across benchmarks and reflecting stronger generalization to diverse tasks. These findings highlight that systematically harnessing failures as structured learning signals, rather than treating them as byproducts, opens a promising direction for advancing self-improving agents.
@@
  • We introduce a failure agent that, unlike prior frozen negative agents trained on human-curated data, continuously learns from failure trajectories and captures a fine-grained failure landscape.
  • We propose a co-evolving agents framework where a target and failure agent improve jointly, with the failure agent generating hard negatives that sharpen decision boundaries and enhance generalization.
  • Our experiments further confirm that failures, when systematically harnessed, can be transformed into structured learning signals that drive more robust self-improving agents.
@@ 2 Related work
Building high-performing agents requires high-quality datasets, which are costly and often infeasible in real-world scenarios. Self-improving agents address this by autonomously generating, refining, and reusing data for continual learning. Some approaches synthesize trajectories from tutorials, documentation, or persona hubs (learnbyinteract; agenttuning; agentrefine), while others use planning methods such as Monte Carlo Tree Search (MCTS) (agentR). Beyond dialogue, self-improvement has been explored in programmatic action composition (dynasaur), robotics (robocat), and code generation (golden). Another line leverages failure trajectories paired with expert ones for preference optimization (eto; IPR), but these typically use failures as-is, limiting generalization. Multi-agent variants (macpo) employ negative agents trained on curated failure datasets, yet these are frozen and restricted to dialogue tasks, offering only limited benefit compared to success-based supervision.
@@
Reinforcement Learning from Human Feedback (RLHF) (lee2024rlaif) has been the standard paradigm for aligning language models, but it requires costly reward modeling and policy optimization. Recent contrastive methods such as Direct Preference Optimization (DPO) (dpo) and Generalized Preference Optimization (GRPO) (grpo) simplify this process by directly optimizing policies on preference pairs, bypassing explicit reward models. At the core of contrastive optimization is the idea that learning benefits most from informative comparisons. In particular, hard negatives that are difficult to distinguish from the preferred ones and thus yield small preference margins, are known to provide stronger supervision and promote sharper decision boundaries (hardnegative; dpo; simclr).
@@ 3 Preliminaries
The interaction between an LLM agent and its environment can be formalized as a partially observable Markov decision process (POMDP)  $(\mathcal{U},\mathcal{S},\mathcal{A},\mathcal{O},T,R)$, as in eto. Here, $\mathcal{U}$ denotes the instruction space, $\mathcal{S}$ the state space, $\mathcal{A}$ the action space, $\mathcal{O}$ the observation space, $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ the transition function, and $R:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ the reward function. In our LLM-agents setting, $\mathcal{U},\mathcal{A},$ and $\mathcal{O}$ are expressed in natural language.
@@
$\pi_{\theta}(e\mid u)=\prod_{j=1}^{n}\pi_{\theta}(a_{j}\mid u,a_{1},o_{1},\ldots,o_{j-1}),$ (2)
@@ 4 Method
In this section, we propose our co-evolving agents framework in which a target agent and a failure agent are trained in alternating phases and gradually improve through mutual interaction. In Section 4.1, we describe the behavioral cloning stage used to initialize the base policy from supervised fine-tuning on expert trajectories. Next, in Section 4.2, we introduce the failure agent, which learns via preference optimization over failure trajectories from both the target and itself and generates fine-grained hard negatives that are close to success yet still failures. Finally, in Section 4.3, we describe how the target agent leverages expert trajectories, its own predicted failures, and failures generated by the failure agent to construct diverse preference datasets for direct preference optimization (DPO) (dpo). By training on this richer set of comparisons and alternating with the failure agent, the target agent achieves more effective learning and stronger generalization within a co-evolutionary loop. The overall pipeline is illustrated in Figure 1.
@@ 4.1 Behavioral Cloning with Supervised Fine-tuning
$\mathcal{L}_{\mathrm{SFT}}(\theta)=-\mathbb{E}_{e\sim\mathcal{D}}\big[\log\pi_{\theta}(e\mid u)\big],$ (3)
@@
$\pi_{\theta}(e\mid u)=\prod_{t=1}^{n}\pi_{\theta}(a_{t}\mid u,a_{<t},o_{<t}).$ (4)
@@
$\mathcal{L}_{\mathrm{SFT}}(\theta)=-\sum_{k=1}^{l}\log\pi_{\theta}(t_{k}\mid t_{<k})\cdot\mathbf{1}(t_{k}\in\mathcal{A}),$ (5)
@@
This supervised fine-tuning stage provides the base policy $\pi_{\mathrm{base}}$, which serves as the starting point for co-evolution with the failure agent. To ensure simplicity, both target and failure agents are initialized from independently trained base policies on the same expert dataset, providing a comparable starting point while allowing only minor stochastic differences.
@@ 4.2 Failure Agent for Generating Hard Negatives
We introduce an auxiliary failure agent $\pi_{\theta_{f}}$ whose role is to specialize in unsuccessful trajectories and refine them into informative hard negatives . Unlike the target agent $\pi_{\theta_{t}}$, which is optimized toward expert success, the failure agent focuses on modeling the space of failures and extracting fine-grained signals from them. This complementary specialization enables the two agents to co-evolve in alternating phases.
@@
The preference dataset for the failure agent consists of failure trajectories generated by both the target and itself. Formally, let $\mathcal{F}_{\text{tgt}}=\{e_{\text{tgt}}\mid r(u,e_{\text{tgt}})<1\}$ and $\mathcal{F}_{\text{fail}}=\{e_{\text{fail}}\mid r(u,e_{\text{fail}})<1\}$ denote the sets of failure trajectories generated by the target and failure agents, respectively. We construct a preference dataset by pairing failures with different reward levels:
$\mathcal{D}_{\text{fail}}=\big\{(u,e_{\text{chosen}},e_{\text{rejected}})\,\big|\,e_{\text{chosen}},e_{\text{rejected}}\in\mathcal{F}_{\text{tgt}}\times\mathcal{F}_{\text{fail}},\;r(u,e_{\text{chosen}})>r(u,e_{\text{rejected}})\big\}.$ (6)
Here, both $e_{\text{chosen}}$ and $e_{\text{rejected}}$ are failure trajectories, and their relative preference is determined by evaluating rewards with the given reward model, where the higher-reward trajectory is selected as $e_{\text{chosen}}$ and the lower as $e_{\text{rejected}}$. This construction enables the failure agent to exploit not only its own generated failures but also those produced by the target, providing a richer set of pairwise comparisons.
@@
We adopt the direct preference optimization (DPO) objective (dpo) for training on failure trajectories. Given a reference policy $\pi_{\text{ref}}$, the failure agent $\pi_{\theta_{f}}$ is updated by
$\mathcal{L}_{\text{DPO}}(\theta_{f})=-\mathbb{E}_{(u,e_{\text{chosen}},e_{\text{rejected}})\sim\mathcal{D}_{\text{fail}}}\Big[\log\sigma\big(\beta\log\tfrac{\pi_{\theta_{f}}(e_{\text{chosen}}\mid u)}{\pi_{\text{ref}}(e_{\text{chosen}}\mid u)}-\beta\log\tfrac{\pi_{\theta_{f}}(e_{\text{rejected}}\mid u)}{\pi_{\text{ref}}(e_{\text{rejected}}\mid u)}\big)\Big],$ (7)
where $u$ is the task instruction, $\sigma(\cdot)$ is the logistic sigmoid, and $\beta$ is a scaling factor. This objective drives the failure agent to distinguish between relatively better and worse failures, thereby capturing subtle distinctions within the failure space.
@@
By constructing preference datasets over pairs of failure trajectories and training with the DPO objective, the failure agent learns to capture nuanced differences among failures rather than only aligning with expert trajectories. This enables fine-grained learning of the failure trajectory landscape by considering diverse failure cases and, in particular, generating near-success failures that remain informative despite not solving the task. Such hard negatives provide informative signals that cannot be obtained from simple expert-versus-failure comparisons, allowing the model to establish a sharper decision boundary between success and failure. Please note that we define hardness based on the relative reward value rather than using a fixed threshold. Within the 0–1 reward range where success corresponds to 1.0, a failure trajectory with a reward of 0.8 is considered harder than one with 0.5 because it lies closer to the success reward. Incorporating these refined failure trajectories into the target agent’s optimization enables the model to learn sharper decision boundaries and achieve stronger generalization. To better understand the role of the failure agent, we further conduct both quantitative and qualitative analyses of the generated failure trajectories (Section 5.2).

### v4

@@ Abstract
Recent advances in large foundation models have enabled task-specialized agents, yet their performance remains bounded by the scarcity and cost of high-quality training data. Self-improving agents mitigate this bottleneck by using their own trajectories in preference optimization, which pairs these trajectories with limited ground-truth supervision. However, their heavy reliance on self-predicted trajectories prevents the agent from learning a coherent preference landscape, causing it to overfit to the scarce supervision and resulting in limited improvements. To address this, we propose a co-evolving agents framework in which an auxiliary failure agent generates informative hard negatives and co-evolves with the target agent by learning from each other’s failure trajectories. The failure agent learns through preference optimization using only failure trajectories from both agents, thereby generating hard negatives that are close to success yet remain failures. Incorporating these informative hard negatives into the target agent’s preference optimization refines the preference landscape and improves generalization. Our comprehensive experiments on complex multi-turn benchmarks such as web shopping, scientific reasoning, and SQL tasks show that our framework generates higher-quality hard negatives, leading to consistent improvements over baselines. These results demonstrate that failures, rather than being used as-is, can be systematically transformed into structured and valuable learning signals in self-improving agents.
@@ 1 Introduction
The rapid progress of large foundation models (OpenAI, 2025; Yang et al., 2025; MetaAI, 2025; Anthropic, 2025; Gemini Team, 2025) has facilitated the rise of task-specialized agents across diverse domains, from open-domain dialogue to scientific reasoning (SU et al., 2025; Zeng et al., 2024; Fu et al., 2025; Bousmalis et al., 2024). These agents inherit the broad generalization capacity of pretrained models, enabling effective adaptation to new tasks with relatively limited supervision. This capability has motivated growing interest in adapting foundation models into reliable and effective domain-specialized agents. Recent advances in multi-agent systems and preference optimization further highlight the potential of combining broad pretraining with specialized adaptation.
@@
Nevertheless, the effectiveness of such agents remains constrained by the quality of task-specific training data (Zhou et al., 2024; Zhao et al., 2024). High-quality datasets are essential for reliable reasoning and decision making, providing the signals required for adaptation to specialized domains. Yet, constructing such datasets is expensive and labor-intensive, often requiring domain expertise and extensive annotation. In many real-world scenarios, building large curated datasets is infeasible, and the need to repeatedly curate data to keep pace with non-stationary environments makes this approach impractical. This bottleneck has motivated growing interest in methods that enable agents to improve autonomously without relying on continuous manual dataset curation (Yuan et al., 2025b; Nguyen et al., 2025; Yin et al., 2025). A promising direction is to automatically curate training signals from agent interactions, allowing learning to scale beyond static, human-labeled corpora. The central challenge is to transform abundant but noisy interaction data into structured supervision that drives reliable improvement.
@@
Self-improving agents (SU et al., 2025; Zeng et al., 2024; Fu et al., 2025) have emerged as a promising paradigm to reduce reliance on costly human annotation. Maintaining agents at state-of-the-art performance would require continuous human annotation, which is prohibitively costly and infeasible at scale. Instead, self-improving agents automate parts of the data construction process by synthesizing expert-like trajectories from external resources such as documentation or databases, and by repurposing predicted failures as preference data for training. However, in challenging downstream tasks where pretrained LLMs perform poorly, generating high-quality trajectories themselves remains a major bottleneck, making fully autonomous self-improvement difficult in practice.
@@
In this context, Exploration-Based Trajectory Optimization (ETO) (Song et al., 2024) constructs preference datasets by pairing the agent’s own predicted failure trajectories with ground-truth ones using a given reward model, enabling preference optimization (Rafailov et al., 2023). Despite the promise of autonomous improvement, this approach remains limited, as it relies on the model’s raw failure trajectories. These failures are substantially less informative than human-curated negatives, especially in tasks where pretrained LLMs lack prior knowledge. As a result, the dispreferred trajectories offer only weak contrast, making DPO focus on simply increasing the likelihood of expert trajectories rather than shaping a fine-grained preference landscape. This weak supervision causes the model to gravitate toward the small set of expert trajectories, ultimately leading to overfitting rather than robust task understanding.
@@
To address this, we propose a co-evolving agents framework in which an auxiliary failure agent generates informative hard negatives and co-evolves with the target agent by learning from each other’s failure trajectories. The failure agent learns through preference optimization using only failure trajectories from both agents, thereby generating hard negatives that are close to success yet remain failures. Incorporating these informative hard negatives into the target agent’s preference optimization refines the preference landscape and improves generalization. Crucially, this design enables the agent to autonomously generate informative hard negatives (Robinson et al., 2021; Rafailov et al., 2023; Chen et al., 2020) that are high-reward failures close to success, without any external supervision. Incorporating these hard negatives into the target agent’s preference optimization provides stronger and more diverse contrastive signals, refining the preference landscape and improving generalization.
We validate our framework through comprehensive analysis and experiments across diverse domains. These evaluations are conducted on complex multi-turn benchmarks, including the online shopping environment WebShop (Yao et al., 2022), the science reasoning environment ScienceWorld (Wang et al., 2022), and the interactive SQL environment InterCodeSQL (Yang et al., 2023). Our quantitative and qualitative analysis of failure trajectories on these benchmarks shows that the proposed failure agent does not merely imitate expert trajectories but consistently generates high-reward failures that serve as informative hard negatives. Experiments across a diverse models further demonstrate substantial improvements over competitive baselines, achieving large gains on all benchmarks and reflecting stronger generalization to diverse tasks. These findings highlight that systematically harnessing failures as structured learning signals, rather than treating them as byproducts, opens a promising direction for advancing self-improving agents.
@@
  • We introduce a failure agent that continuously learns from failure trajectories, modeling a fine-grained failure landscape and autonomously generating high-reward failures as informative hard negatives.
  • We propose a co-evolving agents framework where the target and failure agents improve jointly, with the failure agent generating hard negatives that strengthen preference optimization and improve generalization.
  • Through quantitative and qualitative experiments on complex multi-turn benchmarks, we show that systematically harnessed failures become structured learning signals that yield consistent gains.
@@ 2 Related work
Building high-performing agents requires high-quality datasets, which are costly and often infeasible in real-world scenarios. Self-improving agents address this by autonomously generating, refining, and reusing data for continual learning. Some approaches synthesize trajectories from tutorials, documentation, or persona hubs (SU et al., 2025; Zeng et al., 2024; Fu et al., 2025), while others use planning methods such as Monte Carlo Tree Search (MCTS) (Yuan et al., 2025b). Beyond dialogue, self-improvement has been explored in programmatic action composition (Nguyen et al., 2025), robotics (Bousmalis et al., 2024), and code generation (Yin et al., 2025). Another line leverages failure trajectories paired with expert ones for preference optimization (Song et al., 2024; Xiong et al., 2024), but these typically use failures as-is, limiting generalization. Multi-agent variants (Zhang et al., 2024) employ negative agents trained on curated failure datasets, yet these are frozen and restricted to dialogue tasks, offering only limited benefit compared to success-based supervision.
@@
Reinforcement Learning from Human Feedback (RLHF) (Lee et al., 2024) has been the standard paradigm for aligning language models, but it requires costly reward modeling and policy optimization. Recent contrastive methods such as Direct Preference Optimization (DPO) (Rafailov et al., 2023) and Generalized Preference Optimization (GRPO) (Tang et al., 2024) simplify this process by directly optimizing policies on preference pairs, bypassing explicit reward models. At the core of contrastive optimization is the idea that learning benefits most from informative comparisons. In particular, hard negatives that are difficult to distinguish from the preferred ones and thus yield small preference margins, are known to provide stronger supervision and promote sharper decision boundaries (Robinson et al., 2021; Rafailov et al., 2023; Chen et al., 2020).
@@ 3 Preliminaries
The interaction between an LLM agent and its environment can be formalized as a partially observable Markov decision process $(\mathcal{U},\mathcal{S},\mathcal{A},\mathcal{O},T,R)$, as in Song et al. (2024). Here, $\mathcal{U}$ denotes the instruction space, $\mathcal{S}$ the state space, $\mathcal{A}$ the action space, $\mathcal{O}$ the observation space, $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ the transition function, and $R:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ the reward function. In our LLM-agents setting, $\mathcal{U},\mathcal{A},$ and $\mathcal{O}$ are expressed in natural language.
@@
$\pi_{\theta}(e\mid u)=\prod_{j=1}^{n}\pi_{\theta}(a_{j}\mid u,a_{1},o_{1},\ldots,o_{j-1}),$
@@ 4 Method
In this section, we present our co-evolving agents framework, where a target agent and a failure agent improve jointly through alternating training phases. Section 4.1 outlines the behavioral cloning stage for initializing the target policy, and Section 4.2 introduces the failure agent, which learns from failure trajectories and generates fine-grained hard negatives. Section 4.3 then describes how the target agent learns from the hard negatives generated by the failure agent through DPO (Rafailov et al., 2023), completing the co-evolutionary loop. The overall pipeline is illustrated in Figure 1.
@@ 4.1 Behavioral Cloning with Supervised Fine-tuning
$\mathcal{L}_{\mathrm{SFT}}(\theta)=-\mathbb{E}_{e\sim\mathcal{D}}\big[\log\pi_{\theta}(e\mid u)\big],$
@@
$\pi_{\theta}(e\mid u)=\prod_{t=1}^{n}\pi_{\theta}(a_{t}\mid u,a_{<t},o_{<t}).$
@@
$\mathcal{L}_{\mathrm{SFT}}(\theta)=-\sum_{k=1}^{l}\log\pi_{\theta}(t_{k}\mid t_{<k})\cdot\mathbf{1}(t_{k}\in\mathcal{A}),$
@@
This supervised fine-tuning stage provides the base policy $\pi_{\mathrm{base}}$, which serves as the starting point for both the target and failure agents. Using the same base policy for both agents ensures a comparable starting point while allowing only minor stochastic differences during training.
@@ 4.2 Failure Agent for Generating Hard Negatives
We introduce an auxiliary failure agent $\pi_{\theta_{f}}$ that specializes in learning from failure trajectories and transforming them into fine-grained, informative hard negatives. Unlike the target agent $\pi_{\theta_{t}}$, which is optimized toward expert success, the failure agent focuses solely on modeling the failure landscape. This complementary specialization enables the two agents to co-evolve through alternating training phases.
@@
The preference dataset for the failure agent consists of failure trajectories generated by both the target and itself. Formally, let $e_{\text{tgt}}$ and $e_{\text{fail}}$ denote trajectories generated by the target agent and the failure agent, respectively. We define the corresponding failure sets as $\mathcal{F}_{\text{tgt}}=\{e_{\text{tgt}}\mid r(u,e_{\text{tgt}})<1\}$ and $\mathcal{F}_{\text{fail}}=\{e_{\text{fail}}\mid r(u,e_{\text{fail}})<1\}$ denote the sets of failure trajectories generated by the target and failure agents, respectively. We construct a preference dataset by pairing failures with different reward levels:
$\mathcal{D}_{\text{fail}}=\Big\{(u,e^{+},e^{-})\,\Big|$ $e^{+},e^{-}\in\mathcal{F}_{\text{tgt}}\times\mathcal{F}_{\text{fail}}\Big\}.$
Here, $e^{+}$ denotes the preferred trajectory and $e^{-}$ the dispreferred one, with preference determined by their reward values. In this setting, both $e^{+}$ and $e^{-}$ are failure trajectories, where the higher-reward failure is assigned to $e^{+}$ and the lower-reward one to $e^{-}$. This pairing allows the failure agent to leverage failures from both agents, yielding a richer set of comparisons. In addition, the target agent’s failures further provide weak supervision that guides the failure agent toward competent behaviors and prevents collapse into trivial failures.
@@
We adopt the direct preference optimization (DPO) objective (Rafailov et al., 2023) for training on failure trajectories. Given a reference policy $\pi_{\text{ref}}$, the failure agent $\pi_{\theta_{f}}$ is updated by
$\mathcal{L}_{\text{DPO}}(\theta_{f})=-\mathbb{E}_{(u,e^{+},e^{-})\sim\mathcal{D}_{\text{fail}}}\Big[\log\sigma\Big(\beta\log\tfrac{\pi_{\theta_{f}}(e^{+}\mid u)}{\pi_{\text{ref}}(e^{+}\mid u)}$
$\qquad\qquad-\beta\log\tfrac{\pi_{\theta_{f}}(e^{-}\mid u)}{\pi_{\text{ref}}(e^{-}\mid u)}\Big)\Big].$
where $u$ is the task instruction, $\sigma(\cdot)$ is the logistic sigmoid, and $\beta$ is a scaling factor. This objective encourages the failure agent to learn fine-grained distinctions within the failure landscape, enabling it to identify both diverse and high-reward failures as hard negatives.
@@
Preference optimization depends on strong contrast between preferred and dispreferred trajectories, yet weak or distant negatives often yield trivial comparisons that cause the target agent to collapse into overly simple modes with poor generalization. Our failure agent overcomes this limitation by learning fine-grained distinctions across diverse failure trajectories and generating near-success failures that remain informative despite not solving the task. These hard negatives provide contrastive signals that simple expert-versus-failure pairs cannot offer, and incorporating them into the target agent’s DPO training sharpens the preference landscape and leads to substantially stronger generalization. To better understand the role of the failure agent, we further conduct both quantitative and qualitative analyses of the generated failure trajectories (Section 5.2).